OpenStudy (anonymous):

Solve x - 1/11 = 2/3

OpenStudy (swag):

X=25/33

OpenStudy (anonymous):

Assume that the expression is: (x - 1)/11 = 2/3 Cross-multiply.. 3(x - 1) = 11 * 2 3x - 3 = 22 Then, rearrange the terms.. 3x - 3 + 3 = 22 + 3 3x = 25 x = 25/3 I hope this helps!